1 -1 1 1. X= 0 1 3 X 4 3 1Solve using linear Differential equations
Answer #1

2=52644) 21261). aucto Solution :- for a system of fisst andler defferential equations. AX, LE* Suppose that A has in eigen11 = 4 Consicles the System (A=9, IX=0 sou nt 2- 263 E3-1 24 0-33 le 43-3 913 -571-77127213-0-YO - 3212+3713-0 All +372-32637> 0 -N-N2 +213-0 -262 +326=0 404+5=113=0 2 1/2 = 5013 @=> = n4=-2013 -11-3763+73=0 1-2013 lou 12 H 262 3262 012 E = 1/₂ = Can=> // 1-7 Can be taken as-an-eigen Veetos of 4 COBASPaneling to 13z-5 Hence the general Solution to the System -1 os X IX is